OpenStudy (anonymous):

How do I simplify this equation?!?! I have tried to several times, but I can't seem to ever get the correct answer. 4 + 2i / (2/3 + 1/2i) Help appreciated ASAP. Thank you. :)

ganeshie8 (ganeshie8):

guess i wud first get rid of fractions

ganeshie8 (ganeshie8):

multiply 6 top and bottom

OpenStudy (anonymous):

May I ask why 6?

OpenStudy (shamil98):

2/3 x 6 = 12/3 = 4 1/2 x 6 = 6/2 = 3 to get whole numbers

ganeshie8 (ganeshie8):

\(\huge \frac {4 + 2i}{\frac{2}{3} + \frac{1}{2}i}\)

ganeshie8 (ganeshie8):

^ is ur expression like above ?

OpenStudy (anonymous):

yes that is the equation

ganeshie8 (ganeshie8):

shamil explained nicely why 6 did u get why 6 ? :)

OpenStudy (anonymous):

yes i do now

ganeshie8 (ganeshie8):

\(\huge \frac {4 + 2i}{\frac{2}{3} + \frac{1}{2}i} \) \(\huge \frac {6 \times (4 + 2i)}{6 \times (\frac{2}{3} + \frac{1}{2}i)} \)

OpenStudy (anonymous):

I now have 24 + 12i / 4 + 31 .

OpenStudy (shamil98):

\[\frac{ 24 + 12i }{ 4 + 3i }\]

OpenStudy (anonymous):

how will that further simplify? Would I find the conjugate of the denominator then multiply the top and bottom by it?

ganeshie8 (ganeshie8):

\(\huge \frac {4 + 2i}{\frac{2}{3} + \frac{1}{2}i} \) \(\huge \frac {6 \times (4 + 2i)}{6 \times (\frac{2}{3} + \frac{1}{2}i)} \) \(\huge \frac {24+12i}{4+3i} \)

ganeshie8 (ganeshie8):

exactly !

ganeshie8 (ganeshie8):

multiply top and bottom wid conjugate of denominator, so that u get the complex number in standard form

OpenStudy (anonymous):

I got 132 - 24i / 25 , but that is not one of my choices. Did I do something wrong?

OpenStudy (anonymous):

Could 132 - 24i / 25 become (132 / 25) - (24i / 25) ?

OpenStudy (shamil98):

Yes.

OpenStudy (shamil98):

You got the answer correct, but it can be written differently.

OpenStudy (shamil98):

It's like 4 - 3 / 5 = 4/5 - 3/5

OpenStudy (anonymous):

Thank you all so very much!! :)